OpenStudy (anonymous):

what is et=689 (solve for t)exponential function

OpenStudy (mertsj):

\[e ^{t}= 689\]

OpenStudy (anonymous):

t=ln(689)=6.53524

OpenStudy (anonymous):

bc ln(e^t)=t

OpenStudy (mertsj):

t(ln(e)) =ln689 Since the ln of e = 1, t = ln(689)